# Accelerating the Search for Superconductors Using Machine Learning

### v1

@@ Abstract
Prediction of critical temperature ($T_{c}$) of a superconductor remains a significant challenge in condensed matter physics. While the BCS theory explains superconductivity in conventional superconductors, there is no framework to predict critical temperature of unconventional, higher $T_{c}$ superconductors. Quantum Structure Diagrams (QSD) were successful in establishing structure-property relationship for superconductors, quasicrystals, and ferroelectric materials starting from chemical composition. Building on the QSD ideas, we first demonstrate that the principal component analysis of superconductivity data naturally uncovers the clustering of various classes of superconductors. We use these ideas along with machine learning analysis and intelligently cleaned and merged databases of superconductors to develop predictive models of critical temperature of a superconductor using its chemical composition and attributes. Earlier studies generally relied on datasets with inconsistencies and incomplete information, leading to suboptimal predictions. To address this, we introduce a data-cleaning workflow to enhance the statistical quality of superconducting material databases by eliminating redundancies and resolving inconsistencies. With this improvised database, we apply a supervised machine learning framework and develop a Random Forest model to predict superconductivity and critical temperature as a function of descriptors motivated from Quantum Structure Diagrams. We demonstrate that this model generalizes effectively in reasonably accurate prediction of $T_{c}$ of compounds outside the database. We further employ our model to systematically screen materials across multiple materials databases as well as various chemically plausible combinations of elements and predict $\mathrm{Tl}_{5}\mathrm{Ba}_{6}\mathrm{Ca}_{6}\mathrm{Cu}_{9}\mathrm{O}_{29}$ to exhibit superconductivity with a $T_{c}$ $\sim$ 105 K. Being based on the descriptors used in QSD’s, our model bypasses structural information and predicts $T_{c}$ merely from the chemical composition, and hence is quite efficient and powerful in the search for new superconductors.
@@ I Introduction
Superconductors are broadly of two types: Conventional and Unconventional superconductors. Conventional superconductors are those where the mechanism of superconductivity is based on electron-phonon coupling within the BCS Theory [11], where phonons provide attractive interactions between two electrons that form a Cooper pair. Compounds like intercalated graphite (e.g. $\text{C}_{6}\text{Yb}$ [12] ), A15 compounds (e.g. $\text{V}_{3}\text{Si}$ [3] ), elemental compounds (e.g. Hg [1] ), and metal hydrides (e.g $\text{CeH}_{9}$ [13] ) belong to this class. Eliashberg Theory, an extension of the BCS theory that incorporates the full electron-phonon interaction [14], complemented by the empirical equation proposed by McMillan [15] and Allen-Dynes [16] estimates the critical temperature of superconductors based on electron-phonon coupling and average phonon frequency.
@@
This formalism can be readily used within the framework of first-principles Density Functional Theory to predict critical temperature [17] of conventional superconductors. Among the known conventional superconductors at ambient pressure, the highest critical temperature is observed in Magnesium Diboride ($\textnormal{MgB}_{2}$) with a critical temperature of 39 K. [18].
@@
High temperature superconductors, like cuprates, iron-based compounds, are unconventional superconductors, where there are various mechanisms like hole theory [19], resonating valence bond theory [20], however not with quantitative prediction of critical temperature. While achieving superconductivity at room temperature is a dream of condensed matter physicists, predictive model of critical temperature is a serious challenge. The high computational costs for screening thousands of compounds using Density Functional Theory (DFT) makes it resource intensive. Additionally, DFT based methods are limited to predict critical temperature in a weakly correlated conventional systems, thereby hindering the exploration for materials with potential for room temperature superconductivity.
@@
Statistical methods have been powerful tools, even before the conceptualization of machine learning. One such statistical approach, developed in the 1980’s by J.C. Phillips involved using Quantum Structural Diagrams [21] to establish structure-property relationships based on chemical composition.
@@
Many attempts have been made since 2018 to estimate critical temperature using machine learning. Stanev et al.[22] used 145 MAGPIE descriptors and 26 AFLOW features based on elemental composition to predict critical temperature using Random Forest method. Hamidieh’s [23] work presented a data-driven statistical model to predict critical temperature of superconductors using only the chemical formulas of the materials. The model extracted 81 features based on the elemental properties. Gashmard et al.[24] used 322 atomic features generated from Soraya package in Python to predict critical temperature of superconductors. Zhang et al.[25] used machine learning with structural descriptors and electronic properties from the AFLOW database to predict the critical temperature of superconductors. Jung et al.[26] developed a framework that uses gradient-boosting and statistical feature selection to predict critical temperature. Konno et al.[27] introduced a novel deep learning model that learns to interpret the periodic table, rather than relying on manually engineered features. The model was trained on the materials from SuperCon and Crystallographic Open Database (COD). Kaplan et al.[28] employed a convolutional neural network (CNN) to analyze the chemical compositions of materials without augmenting non-superconductors in their model. Pereti et al.[29] utilized a Deep Set neural network that processes chemical compositions as sets, thus eliminating the bias from the ordering of the elements. The model was trained on data from the SuperCon and Crystallographic Open Database (COD) databases.
@@
Choudhary et al.[30] presented the j-Scr workflow, which combines DFT-based screening with empirical rules and deep learning to predict superconductor transition temperatures. The workflow relies on the McMillan-Allen-Dynes formula. Cerqueira et al.[31] combined ab-initio calculations of electron-phonon coupling using Density-Functional Perturbation Theory (DFPT) with machine learning to predict critical temperature in conventional superconductors. The use of the machine learning model MODnet allowed them to reduce the number of calculations by focusing on the promising candidates in the material search. Seegmiller et al.[32] introduced the DiSCoVeR algorithm and emphasized the importance of a composition-based approach for identifying new superconductors.
@@
In this work, we focus on utilizing a cleaner database of superconducting materials along with chemical composition-based descriptors generated from Quantum Structure Diagrams, to predict superconductivity and critical temperature using classical machine learning algorithms.
@@
In this paper, in Sec II, we discuss on features and descriptor generation from chemical composition using Quantum Structure Diagrams [21] to be used to train a machine learning model. In Sec III, we describe our data cleaning workflow for preparing the SuperCon-MTG dataset. Section IV presents an analysis of the data distribution in our modified database, a Principal Component Analysis (PCA) study of superconductors, and the use of a Random Forest model for classification and regression tasks. We also perform SHAP [33] analysis to interpret the role of each feature in predicting the critical temperature, identify key features, and train a model using only these features. Finally, in Sec. V, we demonstrate the search for new superconducting materials by validating our model’s performance on recently discovered compounds and exploring potential superconductors in the Materials Project [34] database.
@@ II Descriptors from QSD’s and Engineered features
Machine learning models rely on features that effectively predict the target variable. Superconducting critical temperature can depend on various factors: In conventional superconductors, factors like electron-phonon coupling constant ($\lambda$), Debye temperature ($\theta_{\textnormal{D}}$) play important roles, whereas in unconventional superconductors factors such as hole doping, Cu-O bond lengths become significant.
Before the advent of the BCS theory in the 1950’s, discovery relied on empirical rules to find potential superconductors. One such rules was given by Matthias [35, 36], who identified over hundred’s of superconducting materials, primarily alloys, by following heuristic principles. Among these, the average valence electron count per atom (Z) emerged as an important predictor, with superconductivity observed only when Z was between 2 and 8. The rule suggested that materials with high symmetry, preferably cubic or hexagonal, are plausible candidates for superconductivity. However, Matthias’ rule failed to account for the absence of superconductivity in metals like Mg, which has Z = 2, and in La-Ba-Cu-O materials which possess a layered structure.
The three golden descriptors of Quantum Structural diagrams are: (a) Weighted average of metallic electronegativity difference ($\overline{\Delta\chi}$), (b) Weighted average of orbital radii difference ($\overline{\Delta R}$) and (c) Weighted average of valence electron number ($\overline{N_{v}}$). It was successful in organizing crystal structure types of binary and ternary alloys [37], as well in identifying new quasicrystals [38]. Villars et al. [39] showed that for 60 known superconductors with $T_{\textnormal{c}}$ $>10\text{K}$, these superconductors were found in one of the three islands in $T_{\textnormal{c}}$ v/s $\overline{N_{v}}$ plane. The first island was dominated by the A15 compounds, the second by the borides, carbides and nitrides and the third one by the Cu-oxide, chevrel sulphide and selenide superconductors. Later, Rabe et al. [40] extended this analysis to 70 known superconductors at that time and showed a similar pattern. Using Quantum Structural Diagrams, they uncovered that oxide ferroelectric materials were localized in a smaller region. This was significant because it was not merely a consequence of specific structural types but rather a reflection of the underlying chemical factors responsible for ferroelectricity.
Inspired by these, we explore if these descriptors can be used with machine learning to predict critical temperature of superconductors.
@@
where ‘$\chi$’ is the electronegativity in Martynov - Batsanov scale, ‘$R$’ is the Zunger’s pseudopotential radii sums, ‘$N$’ is the valence electron number (see Supplementary Information (SI), Sec. I [41]), $\overline{\Delta\chi}$ is the weighted average of metallic electronegativity difference, $\overline{\Delta R}$ is the weighted average of orbital radii difference, $\overline{N_{v}}$ is the weighted average of valence electron number and subscripts denotes the species ‘A’ and ‘B’.
@@
Quantum Structure Diagrams require the analysis of quaternary and higher-order compounds, treating them as pseudo-ternary compounds by grouping similar elements. However, we extend this approach to treat quaternary and higher-order compounds in a manner similar to binary and ternary compounds. When calculating the weighted average of a property difference, each possible pair is weighted by twice the minimum of their normalized compositions. For the weighted average of a property, the normalized composition of each element is used as the weight factor.
@@
For a binary compound with normalized composition ($\text{A}_{x}\text{B}_{y}$), for example, the descriptors are calculated as follows:
@@
Thus, our feature space consists of six descriptors and 24 statistical features related to four different elemental properties. While superconductivity depends on various factors, including preparation methods, applied pressure, magnetic fields. Our database reports the critical temperature at ambient pressure. As these descriptors do not capture pressure dependence of critical temperature, our model targets prediction of critical temperature at ambient conditions.
Our goal is to develop a model for prediction of a new superconductor. As prediction of structure from chemical composition itself is a daunting problem, we use only chemical descriptors here in developing a machine learning model to predict critical temperature.
@@ III.1 Preprocessing and Data Cleaning
A clean database is essential to analysis with any machine learning algorithm [42]. The notion of “Garbage in, garbage out” holds true, if a model is trained on flawed data, its predictions are bound to be unreliable. In materials science research, data availability is often a significant challenge, constrained by small databases.
Here, we use the SuperCon database maintained by NIMS, Japan [43], which was initiated following the discovery of high-$T_{c}$ superconductors. Three versions of this database are currently available, primarily containing information on the chemical formula, $T_{c}$, the reporting journal, structural details, and measurement methods such as $T_{c}$ onset, the temperature at which the superconducting transition begins, $T_{c}$ at zero resistance, the temperature at which zero resistance is achieved, and magnetization studies for some compounds. The SuperCon database cleaned and published by Stanev et al. [22] in 2018 includes 16,414 compounds. A later version, SuperCon - Ver.220808 [44], contains 33,406 compounds, while the most recent release, SuperCon - Ver.240322 [45], includes 26,358 compounds. Both the 2018 [22] and 2024 [45] databases, have no missing $T_{c}$’s. Unlike, SuperCon - Ver.220808[44], which contains missing entries for $T_{c}$.
@@
    • Compounds with incorrect chemical formulas were identified manually. Their formulas were corrected by referring to the cited journal. E.g., $\text{La}_{0.75}\text{Pr}_{0.25}\text{Ba}_{22}\text{O}_{8}\rightarrow\text{La}_{0.75}\text{Pr}_{0.25}\text{Ba}_{2}\text{O}_{8}$ and $_{10}\rightarrow\text{Os}$ Os.
@@
    Repetition Cleaning: Since our feature generation process involves normalization of the composition, different representations of the same compound become indistinguishable. For example $\text{AB}_{3}$ and $\text{A}_{0.25}\text{B}_{0.75}$ are indistinguishable. We identify such entries in the database and average out their critical temperature values. For SuperCon-Stanev database we simply take the averaged critical temperature. For SuperCon - Ver.220808 and SuperCon - Ver.240322 we retain the compound with averaged $T_{\textnormal{c}}$ only if their standard deviation is less than 5 K . In the SuperCon-Stanev database, we do not apply this standard deviation filter, as critical temperatures in this database will later be updated using the most recent data from SuperCon - Ver.240322.
@@
    Database Comparison: Comparison of the cleaned versions of SuperCon - Ver.220808 and SuperCon - Ver.240322, shows that all 12,950 compounds in SuperCon - Ver.220808 are present in SuperCon - Ver.240322. Similarly, on comparing SuperCon-Stanev with that of SuperCon - Ver.240322, 11,930 compounds in SuperCon-Stanev have critical temperature reported in both the database, while the remaining 3,916 compounds in SuperCon-Stanev have reported critical temperature of 0 K, but are absent in the latest database. Manual search of these compounds shows that some are superconductors, while others are not. It seems that the database of Stanev et al. [22] has imputed the missing data, which is not a good approach for ML model, which would learn from wrong data. Hence, we do manual search of such compounds and update their data back to the database as explained in the next step.
@@
      E.g, $\textnormal{LaFeAs}_{0.3}\text{P}_{0.7}\text{O}$ [46], the phase diagram reported in the cited paper identifies it as an antiferromagnet without superconductivity, We thus update its $T_{\textnormal{c}}$ as 0 K.
@@
    • A machine learning model will be biased if it is trained only on superconductors. To the cleaned database which contains about 1,100 non-superconductors, we add 455 more non-superconductors, along with 43 superconductors reported by Hosono et al [47]. Additionally, through our review of the literature, we identified and manually added 55 materials exhibiting both superconducting and non-superconducting behavior that were not present in the original database. Upon addition of these compounds, some repetition arose, we average the critical temperature of such repeated compounds and retain them only if the standard deviation is less than 5 K. This final cleaned database is referred to as ‘SuperCon-MTG’ named after the ‘Materials Theory Group’. SuperCon-MTG is a cleaned and augmented version of the SuperCon database originally developed and maintained by NIMS Japan [43, 45]. The original dataset is available under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.

### v2

@@ Abstract
Prediction of critical temperature ($T_{c}$) of a superconductor remains a significant challenge in condensed matter physics. While the BCS theory explains superconductivity in conventional superconductors, there is no framework to predict critical temperature of unconventional, higher $T_{c}$ superconductors. Quantum Structure Diagrams (QSD) were successful in establishing structure-property relationship for superconductors, quasicrystals, and ferroelectric materials starting from chemical composition. Building on the QSD ideas, we first demonstrate that the principal component analysis of superconductivity data naturally uncovers the clustering of various classes of superconductors. Earlier machine learning studies generally relied on datasets with inconsistencies and incomplete information, leading to a suboptimal predictions of superconductivity. To address this, we introduce a data-cleaning workflow to enhance the statistical quality of superconducting material databases by eliminating redundancies and resolving inconsistencies. With this improvised database, we apply a supervised machine learning framework and develop a Random forest model to predict superconductivity and critical temperature as a function of descriptors motivated from QSD. Our ML model demonstrates robust performance, achieving an $R^{2}$ score of 0.87 and a mean absolute error of 5.6 K. We demonstrate that this model generalizes effectively in reasonably accurate prediction of $T_{c}$ of compounds outside the database. We further employ our model to systematically screen materials across multiple materials databases as well as various chemically plausible combinations of elements and predict $\mathrm{Tl}_{5}\mathrm{Ba}_{6}\mathrm{Ca}_{6}\mathrm{Cu}_{9}\mathrm{O}_{29}$ to be a superconductor with a $T_{c}$ $\sim$ 105 K. Being based on the descriptors used in QSD’s, our model bypasses structural information and predicts $T_{c}$ merely from the chemical composition, and hence is quite efficient and powerful in the search for new superconductors.
@@ I Introduction
Superconductors are broadly of two types: Conventional and Unconventional superconductors. Conventional superconductors are those where the mechanism of superconductivity is based on electron-phonon coupling within the BCS Theory [Bardeen1957], where phonons provide attractive interactions between two electrons that form a Cooper pair. Compounds like intercalated graphite (e.g. $\text{C}_{6}\text{Yb}$ [Weller2005] ), A15 compounds (e.g. $\text{V}_{3}\text{Si}$ [Hardy1953] ), elemental compounds (e.g. Hg [Onnes1911] ), and metal hydrides (e.g $\text{CeH}_{9}$ [Li2019] ) belong to this class. Eliashberg Theory, an extension of the BCS theory that incorporates the full electron-phonon interaction [Eliashberg1960], complemented by the empirical equation proposed by McMillan [McMillan1968] and Allen-Dynes [Allen1975] estimates the critical temperature of superconductors based on electron-phonon coupling and average phonon frequency.
@@
This formalism can be readily used within the framework of first-principles Density Functional Theory to predict critical temperature [COHEN2010] of conventional superconductors. Among the known conventional superconductors at ambient pressure, the highest critical temperature is observed in Magnesium Diboride ($\textnormal{MgB}_{2}$) with a critical temperature of 39 K [Nagamatsu2001].
@@
High temperature superconductors like cuprates, iron-based compounds are unconventional superconductors, where there are various mechanisms like hole theory [Hirsch1989], resonating valence bond theory [Anderson1987]. However, these mechanisms do not quantitatively predict critical temperature. While achieving superconductivity at room temperature is a dream of condensed matter physicists, developing a predictive model of critical temperature remains a serious challenge. The high computational costs for screening thousands of compounds using Density Functional Theory (DFT) makes it resource intensive. Additionally, DFT based methods are limited to predict critical temperature in a weakly correlated conventional systems. This hinders the exploration for materials with potential for room temperature superconductivity.
@@
Statistical methods have been powerful tools, even before the conceptualization of machine learning. One such statistical approach was developed in the 1980’s by J.C. Phillips. It involved using Quantum Structural Diagrams [Phillips1991] to establish structure-property relationships based on chemical composition.
Many attempts have been made since 2018 to estimate critical temperature using machine learning. Stanev et al.[Stanev2018] used 145 Magpie descriptors and 26 Aflow features based on elemental composition to predict critical temperature using Random forest method. Hamidieh’s [Hamidieh2018] work presented a data-driven statistical model to predict critical temperature of superconductors using only the chemical formulas of the materials. This model extracted 81 features based on the elemental properties.
Gashmard et al.[Gashmard2024] used 322 atomic features generated from Soraya package in Python to predict critical temperature of superconductors. Zhang et al.[Zhang2022] used machine learning with structural descriptors and electronic properties from the Aflow database to predict the critical temperature of superconductors. Jung et al.[Jung2024] developed a framework that uses gradient-boosting and statistical feature selection to predict critical temperature. Trezza et al. [Trezza2023] developed a machine learning classification framework to distinguish superconductors above and below temperature thresholds of 15 K (for benchmarking) and 35 K (for mixed-feature optimization). Their model was based on the Matminer descriptors [Ward2018] and incorporated several techniques, including invariant group identification and SHAP for feature reduction. Classification was performed using entropy-based Quasi Equilibrium Grids and Extra Trees. In their later work [Trezza2024], they extended this approach to address sampling bias between specialized databases like SuperCon [SuperCon2011] and general-purpose databases like the Materials Project (excluding SuperCon entries) [Jain2013], thereby enabling validation of $T_{c}$ predictions, even though the latter may include materials not tested for superconductivity.
@@
Konno et al.[Konno2021] introduced a novel deep learning model that learns to interpret the periodic table, rather than relying on manually engineered features. This model was trained on the materials from SuperCon and Crystallographic Open Database (COD). Kaplan et al.[Kaplan2025] employed a convolutional neural network (CNN) to analyze the chemical compositions of superconductors without augmenting non-superconductors in their model. Pereti et al.[Pereti2023] utilized a Deep Set neural network that processes chemical compositions as sets, thus eliminating the bias from the ordering of the elements. The model was trained on data from the SuperCon and Crystallographic Open Database (COD) databases.
@@
Choudhary et al.[Choudhary2022] presented the j-Scr workflow, which combines DFT-based screening with empirical rules and deep learning to predict superconductor transition temperatures. The workflow relies on the McMillan-Allen-Dynes formula. Cerqueira et al.[Cerqueira2023] combined ab-initio calculations of electron-phonon coupling using Density-Functional Perturbation Theory (DFPT) with machine learning to predict critical temperature in conventional superconductors. The use of the machine learning model MODnet allowed them to reduce the number of calculations by focusing on the promising candidates in the material search. Seegmiller et al.[Seegmiller2023] introduced the DiSCoVeR algorithm and emphasized the importance of a composition-based approach for identifying new superconductors.
@@
In this work, we first focus on developing a cleaner dataset of superconducting materials. Secondly, we use chemical composition-based descriptors derived from the Quantum Structure Diagrams [Phillips1991]. We finally use classical machine learning algorithms to predict superconductivity and critical temperature.
@@
In Sec II, we discuss on features and descriptor generation from chemical composition using Quantum Structure Diagrams [Phillips1991] for training a machine learning model. In Sec III, we describe our data cleaning workflow for preparing the SuperCon-MTG dataset. Section IV presents an analysis of the data distribution in our modified database, a Principal Component Analysis (PCA) study of superconductors, and the use of a Random forest model for classification and regression tasks. We present SHAP [SHAP] analysis to interpret the role of each feature in predicting the critical temperature, identify the key features and finally train a model using only these features. In Sec. V, we demonstrate the search for new superconducting materials and validate our model’s performance with recently discovered compounds and exploring potential superconductors in the Materials Project [Jain2013] database.
@@ II Descriptors from QSD’s and Engineered features
Machine learning models rely on features that effectively predict the target variable. Superconducting critical temperature can depend on various factors. In conventional superconductors, factors like electron-phonon coupling constant ($\lambda$), Debye temperature ($\theta_{\textnormal{D}}$) play important roles. In unconventional superconductors factors such as hole doping, Cu-O bond lengths become significant.
Before the advent of BCS theory in the 1950s, the discovery of superconductors relied on empirical rules. One such rule was given by Matthias [Matthias1957, Matthias1970], who identified over a hundred superconducting materials, primarily alloys, by following heuristic principles. Among these, the average valence electron count per atom (Z) emerged as an important predictor. Superconductivity was observed only when Z was between 2 and 8. This rule suggested that materials with high symmetry, preferably cubic or hexagonal, are plausible candidates for superconductivity. However, Matthias’ rule failed to account for the absence of superconductivity in metals like Mg, which has Z = 2, and in La-Ba-Cu-O materials which possess a layered structure.
The three golden descriptors of Quantum Structural diagrams are: (a) Weighted average of metallic electronegativity difference ($\overline{\Delta\chi}$), (b) Weighted average of orbital radii difference ($\overline{\Delta R}$) and (c) Weighted average of valence electron number ($\overline{N_{v}}$). It was successful in organizing crystal structure types of binary and ternary alloys [Villars1987], as well in identifying new quasicrystals [Villars1986]. Villars et al. [Villars1988] showed that for 60 known superconductors with $T_{\textnormal{c}}$ $10\text{K}$, these superconductors were found in one of the three islands in $T_{\textnormal{c}}$ v/s $\overline{N_{v}}$ plane. The first island was dominated by the A15 compounds, the second by the borides, carbides and nitrides and the third one by the Cu-oxide, chevrel sulphide and selenide superconductors. Later, Rabe et al. [Rabe1992] extended this analysis to 70 known superconductors at that time and uncovered a similar pattern. Using Quantum Structural Diagrams, they uncovered that oxide ferroelectric materials were localized in a smaller region. This was significant because it was not merely a consequence of specific structural types but rather a reflection of the underlying chemical factors responsible for ferroelectricity. Furthermore, Balasubramanian et al. [Bala1989] provided a physical rationale for using electronegativity as a meaningful descriptor. They argue that superconductivity results from the unimpeded motion of Cooper pairs, which requires a near-zero net field, a condition that can hypothetically occur in flat lattice potentials. Since electronegativity governs how atoms attract electrons, it can serve as a proxy. Inspired by these, we explore if these descriptors can be used within machine learning models to predict critical temperature of superconductors.
@@
where $\chi$’ is the electronegativity in Martynov - Batsanov scale, ‘$R$’ is the Zunger’s pseudopotential radii sums, ‘$N$’ is the valence electron number (see Supplementary Information (SI), Sec. I [SI]), $\overline{\Delta\chi}$ is the weighted average of metallic electronegativity difference, $\overline{\Delta R}$ is the weighted average of orbital radii difference, $\overline{N_{v}}$ is the weighted average of valence electron number and subscripts denotes the species ‘A’ and ‘B’.
@@
Quantum Structure Diagrams require the analysis of quaternary and higher-order compounds, treating them as pseudo-ternary compounds by grouping similar elements. However, we extend this approach to treat quaternary and higher-order compounds in a manner similar to binary and ternary compounds. When calculating the weighted average of the property difference, each possible pair is weighted by twice the minimum of their normalized compositions. For the weighted average of a property, the normalized composition of each element is used as a weight.
@@
For a binary compound with normalized composition ($\text{A}_{x}\text{B}_{y}$). The descriptors are calculated as follows:
@@
Thus, our feature space consists of six descriptors and 24 statistical features related to four different elemental properties. While superconductivity depends on various factors, including preparation methods, applied pressure and magnetic fields. Our database reports the critical temperature at ambient pressure. As these descriptors do not capture pressure dependence of critical temperature, our model targets prediction of critical temperature at ambient conditions.
Our goal is to develop a model for prediction of a new superconductor. As prediction of structure from chemical composition itself is a daunting problem, we use only composition based descriptors here in training a machine learning model to predict critical temperature.
@@ III.1 Preprocessing and Data Cleaning
A clean database is essential to analysis with any machine learning algorithm [Pankajakshan2017]. The notion of “Garbage in, garbage out” holds true, if a model is trained on flawed data, its predictions are bound to be unreliable. In materials science research, data availability is often a significant challenge, constrained by small databases.
Here, we use the SuperCon database maintained by NIMS, Japan [SuperCon2011], which was initiated following the discovery of high-$T_{c}$ superconductors. Three versions of this database are currently available, primarily containing information on the chemical formula, $T_{c}$, the reporting journal and structural details. The database also includes measurement methods such as $T_{c}$ onset, the temperature at which the superconducting transition begins, $T_{c}$ at zero resistance, the temperature at which zero resistance is achieved, and magnetization studies for some compounds. The SuperCon database cleaned and published by Stanev et al. [Stanev2018] in 2018 includes 16,414 compounds. A later version, SuperCon - Ver.220808 [SuperCon220808], contains 33,406 compounds, while the most recent release, SuperCon - Ver.240322 [SuperCon240322], includes 26,358 compounds. Both the 2018 [Stanev2018] and 2024 [SuperCon240322] databases, have no missing $T_{c}$’s. Unlike, SuperCon - Ver.220808[SuperCon220808], which contains missing entries for $T_{c}$.
@@
    • Compounds with incorrect chemical formulas were identified manually. Their formulas were corrected by referring to the cited journal. E.g., $\text{La}_{0.75}\text{Pr}_{0.25}\text{Ba}_{22}\text{O}_{8}\rightarrow\text{La}_{0.75}\text{Pr}_{0.25}\text{Ba}_{2}\text{O}_{8}$ and $\text{Os}_{10}\rightarrow\text{Os}$.
@@
    Repetition Cleaning: Since our feature generation process involves normalization of the composition, different representations of the same compound become indistinguishable. For example $\text{AB}_{3}$ and $\text{A}_{0.25}\text{B}_{0.75}$ are indistinguishable. We identify such entries in the database and average out their critical temperature values. For SuperCon-Stanev database we simply take the averaged critical temperature. For SuperCon - Ver.220808 and SuperCon - Ver.240322 we retain the compound with averaged $T_{\textnormal{c}}$ only if their standard deviation is less than 5 K (see Supplementary Information Sec. II [SI] for details). In the SuperCon-Stanev database, we do not apply this standard deviation filter, as critical temperatures in this database will later be updated using the most recent data from SuperCon - Ver.240322.
@@
    Database Comparison: Comparison of the cleaned versions of SuperCon - Ver.220808 and SuperCon - Ver.240322, shows that all 12,950 compounds in SuperCon - Ver.220808 are present in SuperCon - Ver.240322. Similarly, on comparing SuperCon-Stanev with that of SuperCon - Ver.240322, 11,930 compounds in SuperCon-Stanev have critical temperature reported in both the databases. The remaining 3,916 compounds have reported critical temperature of 0 K, but are absent in the latest version of the database. Manual search of these compounds shows that some are superconductors, while others are not. It seems that the database of Stanev et al. [Stanev2018] has imputed the missing data, which is not a good approach for ML model, which would learn from wrong data. Hence, we do manual search of such compounds and update their data back to the database as explained in the next step.
@@
      E.g, $\textnormal{LaFeAs}_{0.3}\text{P}_{0.7}\text{O}$ [Lai2014], the phase diagram reported in the cited paper identifies it as an antiferromagnet without superconductivity, We thus update its $T_{\textnormal{c}}$ as 0 K.
@@
    • A machine learning model will be biased if it is trained only on superconductors. To the cleaned database which contains about 1,100 non-superconductors, we add 455 more non-superconductors, along with 43 superconductors reported by Hosono et al [Hosono2015]. Additionally, through our review of the literature, we identified and manually added 55 materials exhibiting both superconducting and non-superconducting behavior that were not present in the original database. Upon addition of these compounds, some repetition arose, we average the critical temperature of such repeated compounds and retain them only if the standard deviation is less than 5 K. This final cleaned database is referred to as ‘SuperCon-MTG’ named after the ‘Materials Theory Group’. SuperCon-MTG is a cleaned and augmented version of the SuperCon database originally developed and maintained by NIMS Japan [SuperCon2011, SuperCon240322]. The original database is available under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.